$SPAD/src/input unittest2.input

Timothy Daly December 3, 2016

Abstract

Unit test the user level commands

Contents

2

— * —

)set break resume
)sys rm -f unittest2.output
)spool unittest2.output
)lisp (setq *print-circle* t)
)set mes auto off
)clear all
--S 1 of 235
)lisp (identity |$abbreviateTypes|)
--R
--RValue = NIL
--E 1
--S 2 of 235
)lisp (identity |$algebraFormat|)
--R
--RValue = T
--E 2
--S 3 of 235
)lisp (identity |$algebraOutputFile|)
--R
--RValue = "CONSOLE"
--E 3
--S 4 of 235
)lisp (identity |$algebraOutputStream|)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 4
--S 5 of 235 removed this variable
)lisp (identity nil)
--R
--RValue = NIL
--E 5
--S 6 of 235
)lisp (identity |$BreakMode|)
--R
--RValue = |resume|
--E 6
--S 7 of 235
)lisp (identity |$clearExcept|)
--R
--RValue = NIL

--E 7
--S 8 of 235
)lisp (identity |$clearOptions|)
--R
--RValue = (|modes| |operations| |properties| |types| |values|)
--E 8
--S 9 of 235
)lisp (identity |$CommandSynonymAlist|)
--R
--IValue = ((? . "what commands") (|ap| . "what things") ...
--E 9
--S 10 of 235
)lisp (identity |$compileDontDefineFunctions|)
--R
--RValue = T
--E 10
--S 11 of 235
)lisp (identity |$compileRecurrence|)
--R
--RValue = T
--E 11
--S 12 of 235
)lisp (identity compiler::*compile-verbose*)
--R
--RValue = NIL
--E 12
--S 13 of 235
)lisp (identity |$defaultFortranType|)
--R
--RValue = REAL
--E 13
--S 14 of 235
)lisp (identity *default-pathname-defaults*)
--R
--IValue = #p"/tmp/"
--E 14
--S 15 of 235
)lisp (identity |$defaultSpecialCharacters|)
--R
--IValue = (#\^\ #\^[ #\^^#\^_#\O #\- #\\220 #\\255 ...
--E 15

--S 16 of 235
)lisp (identity |$displayDroppedMap|)
--R
--RValue = NIL
--E 16
--S 17 of 235
)lisp (identity |$displayMsgNumber|)
--R
--RValue = NIL
--E 17
--S 18 of 235
)lisp (identity |$displayOptions| )
--R
--IValue = (|abbreviations| |all| |macros| |modes| |names| ...
--E 18
--S 19 of 235
)lisp (identity |$displaySetValue|)
--R
--RValue = NIL
--E 19
--S 20 of 235
)lisp (identity |$displayStartMsgs|)
--R
--RValue = T
--E 20
--S 21 of 235
)lisp (identity |$formulaFormat|)
--R
--RValue = NIL
--E 21
--S 22 of 235
)lisp (identity |$formulaOutputFile|)
--R
--RValue = "CONSOLE"
--E 22
--S 23 of 235
)lisp (identity |$fortIndent|)
--R
--RValue = 6
--E 23
--S 24 of 235
)lisp (identity |$fortInts2Floats|)

--R
--RValue = T
--E 24
--S 25 of 235
)lisp (identity |$fortLength|)
--R
--RValue = 72
--E 25
--S 26 of 235
)lisp (identity |$fortranArrayStartingIndex|)
--R
--RValue = 1
--E 26
--S 27 of 235
)lisp (identity |$fortranDirectory|)
--R
--RValue = "./"
--E 27
--S 28 of 235
)lisp (identity |$fortranFormat|)
--R
--RValue = NIL
--E 28
--S 29 of 235
)lisp (identity |$fortranLibraries|)
--R
--RValue = "-lxlf"
--E 29
--S 30 of 235
)lisp (identity |$fortranOptimizationLevel|)
--R
--RValue = 0
--E 30
--S 31 of 235
)lisp (identity |$fortranOutputFile|)
--R
--RValue = "CONSOLE"
--E 31
--S 32 of 235
)lisp (identity |$fortranPrecision|)
--R
--RValue = |double|

--E 32
--S 33 of 235
)lisp (identity |$fortranSegment|)
--R
--RValue = T
--E 33
--S 34 of 235
)lisp (identity |$fortranTmpDir|)
--R
--RValue = "/tmp/"
--E 34
--S 35 of 235
)lisp (identity |$fortPersistence|)
--R
--RValue = 1
--E 35
--S 36 of 235
)lisp (identity |$fractionDisplayType|)
--R
--RValue = |vertical|
--E 36
--S 37 of 235
)lisp (identity |$frameMessages|)
--R
--RValue = NIL
--E 37
--S 38 of 235
)lisp (identity |$fullScreenSysVars|)
--R
--RValue = NIL
--E 38
--S 39 of 235
)lisp (identity |$giveExposureWarning|)
--R
--RValue = NIL
--E 39
--S 40 of 235
)lisp (identity |$HiFiAccess|)
--R
--RValue = T
--E 40

--S 41 of 235
)lisp (identity |$highlightAllowed|)
--R
--RValue = NIL
--E 41
--S 42 of 235
)lisp (identity |$historyDirectory|)
--R
--RValue = A
--E 42
--S 43 of 235
)lisp (identity |$historyDisplayWidth|)
--R
--RValue = 120
--E 43
--S 44 of 235
)lisp (identity |$historyFileType|)
--R
--RValue = |axh|
--E 44
--S 45 of 235
)lisp (identity |$InitialCommandSynonymAlist|)
--R
--IValue = ((? . "what commands") (|ap| . "what things") ...
--E 45
--S 46 of 235
)lisp (identity |$inputPromptType|)
--R
--RValue = |step|
--E 46
--S 47 of 235
)lisp (identity |$linearFormatScripts|)
--R
--RValue = NIL
--E 47
--S 48 of 235
)lisp (identity $linelength)
--R
--RValue = 77
--E 48
--S 49 of 235
)lisp (identity |$mapSubNameAlist|)

--R
--RValue = NIL
--E 49
--S 50 of 235
)lisp (identity |$mathmlFormat|)
--R
--RValue = NIL
--E 50
--S 51 of 235
)lisp (identity |$mathmlOutputFile|)
--R
--RValue = "CONSOLE"
--E 51
--S 52 of 235
)lisp (identity |$maximumFortranExpressionLength|)
--R
--RValue = 1320
--E 52
--S 53 of 235
)lisp (identity |$nagEnforceDouble|)
--R
--RValue = T
--E 53
--S 54 of 235
)lisp (identity |$nagHost|)
--R
--RValue = "localhost"
--E 54
--S 55 of 235
)lisp (identity |$nagMessages|)
--R
--RValue = T
--E 55
--S 56 of 235
)lisp (identity |$noParseCommands| )
--R
--IValue = (|boot| |copyright| |credits| |fin| |lisp| |pquit| ...
--E 56
--S 57 of 235
)lisp (identity |$oldHistoryFileName|)
--R
--RValue = |last|

--E 57
--S 58 of 235
)lisp (identity |$openMathFormat|)
--R
--RValue = NIL
--E 58
--S 59 of 235
)lisp (identity |$openMathOutputFile|)
--R
--RValue = "CONSOLE"
--E 59
--S 60 of 235
)lisp (identity $openServerIfTrue)
--R
--RValue = T
--E 60
--S 61 of 235
)lisp (identity |$optionAlist|)
--R
--RValue = NIL
--E 61
--S 62 of 235
)lisp (identity |$options|)
--R
--RValue = NIL
--E 62
--S 63 of 235
)lisp (identity |$plainRTspecialCharacters|)
--R
--RValue = (+ + + + |\|| - ? [ ] { } + + + + + |\\|)
--E 63
--S 64 of 235
)lisp (identity |$plainSpecialCharacters0|)
--R
--IValue = (#\N #\N #\N #\N #\O #\` #\o #\\255 #\\275 #\\300 ...
--E 64
--S 65 of 235
)lisp (identity |$plainSpecialCharacters1|)
--R
--IValue = (#\k #\k #\} #\} #\O #\` #\o #\\255 #\\275 #\\300 ...
--E 65

--S 66 of 235
)lisp (identity |$plainSpecialCharacters2|)
--R
--IValue = (#\O #\O #\O #\O #\O #\` #\o #\\255 #\\275 #\\300 ...
--E 66
--S 67 of 235
)lisp (identity |$plainSpecialCharacters3|)
--R
--IValue = (#\` #\` #\` #\` #\O #\` #\o #\\255 #\\275 #\\300 ...
--E 67
--S 68 of 235
)lisp (identity $prettyprint)
--R
--RValue = T
--E 68
--S 69 of 235
)lisp (identity |$printAnyIfTrue|)
--R
--RValue = T
--E 69
--S 70 of 235
)lisp (identity |$printFortranDecs|)
--R
--RValue = T
--E 70
--S 71 of 235
)lisp (identity |$printLoadMsgs|)
--R
--RValue = NIL
--E 71
--S 72 of 235
)lisp (identity |$printMsgsToFile|)
--R
--RValue = NIL
--E 72
--S 73 of 235
)lisp (identity |$printStatisticsSummaryIfTrue|)
--R
--RValue = NIL
--E 73
--S 74 of 235
)lisp (identity |$printTimeIfTrue|)

--S 75 of 235
)lisp (identity |$printTypeIfTrue|)
--R
--RValue = T
--E 75
--S 76 of 235
)lisp (identity |$printVoidIfTrue|)
--R
--RValue = NIL
--E 76
--S 77 of 235
)lisp (identity |$quitCommandType|)
--R
--RValue = |protected|
--E 77
--S 78 of 235
)lisp (identity |$reportBottomUpFlag|)
--R
--RValue = NIL
--E 78
--S 79 of 235
)lisp (identity |$reportCoerceIfTrue|)
--R
--RValue = NIL
--E 79
--S 80 of 235
)lisp (identity |$reportCompilation|)
--R
--RValue = NIL
--E 80
--S 81 of 235
)lisp (identity |$reportEachInstantiation|)
--R
--RValue = NIL
--E 81
--S 82 of 235
)lisp (identity |$reportInstantiations|)
--R
--RValue = NIL

--E 82
--S 83 of 235
)lisp (identity |$reportInterpOnly|)
--R
--RValue = T
--E 83
--S 84 of 235
)lisp (identity |$reportOptimization|)
--R
--RValue = NIL
--E 84
--S 85 of 235
)lisp (identity |$reportSpadTrace|)
--R
--RValue = NIL
--E 85

———-

This generates non-printing characters

--S 86 of 235
--)lisp (identity |$RTspecialCharacters|)
--E 86

— * —

--S 87 of 235
)lisp (identity *standard-input*)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 87
--S 88 of 235
)lisp (identity *standard-output*)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 88
--S 89 of 235
)lisp (identity |$SpadServer|)
--R
--RValue = NIL
--E 89
--S 90 of 235

)lisp (identity $SpadServerName)
--R
--RValue = "/tmp/.d"
--E 90
--S 91 of 235
)lisp (identity |$specialCharacterAlist| )
--R
--IValue = ((|ulc| . 0) (|urc| . 1) (|llc| . 2) (|lrc| . 3) ...
--E 91
--S 92 of 235
)lisp (identity |$specialCharacters|)
--R
--RValue = (+ + + + |\|| - ? [ ] { } + + + + + |\\|)
--E 92
--S 93 of 235
)lisp (identity |$streamCount|)
--R
--RValue = 10
--E 93
--S 94 of 235
)lisp (identity |$streamsShowAll|)
--R
--RValue = NIL
--E 94
--S 95 of 235
)lisp (identity compiler::*suppress-compiler-notes*)
--R
--RValue = T
--E 95
--S 96 of 235
)lisp (identity compiler::*suppress-compiler-warnings*)
--R
--RValue = T
--E 96
--S 97 of 235
)lisp (identity |$systemCommandFunction|)
--R
--RValue = #<compiled-function |InterpExecuteSpadSystemCommand|>
--E 97
--S 98 of 235
)lisp (identity $syscommands)
--R

--IValue = (|abbreviations| |boot| |browse| |cd| |clear| |close| ...
--E 98
--S 99 of 235
)lisp (identity |$systemCommands|)
--R
--IValue = ((|abbreviations| . |compiler|) (|boot| . |development|) ...
--E 99
--S 100 of 235
)lisp (identity *terminal-io*)
--R
--IValue = #<two-way stream 090631d4>
--E 100
--S 101 of 235
)lisp (identity |$testingSystem|)
--R
--RValue = NIL
--E 101
--S 102 of 235
)lisp (identity |$texFormat|)
--R
--RValue = NIL
--E 102
--S 103 of 235
)lisp (identity |$texOutputFile|)
--R
--RValue = "CONSOLE"
--E 103
--S 104 of 235
)lisp (identity |$tokenCommands|)
--R
--IValue = (|abbreviations| |cd| |clear| |close| |compiler| ...
--E 104
--S 105 of 235
)lisp (identity system::*top-level-hook*)
--R
--RValue = RESTART
--E 105
--S 106 of 235
)lisp (identity |$tracedMapSignatures|)
--R
--RValue = NIL
--E 106

--S 107 of 235
)lisp (identity |$traceNoisely|)
--R
--RValue = NIL
--E 107
--S 108 of 235
)lisp (identity |$traceOptionList|)
--R
--IValue = (|after| |before| |break| |cond| |count| |depth| |local| ...
--E 108
--S 109 of 235
)lisp (identity underbar)
--R
--RValue = "_"
--E 109
--S 110 of 235
)lisp (identity |$useEditorForShowOutput|)
--R
--RValue = NIL
--E 110
--S 111 of 235
)lisp (identity |$useFullScreenHelp|)
--R
--RValue = T
--E 111
--S 112 of 235
)lisp (identity |$useInternalHistoryTable|)
--R
--RValue = T
--E 112
--S 113 of 235
)lisp (identity |$useIntrinsicFunctions|)
--R
--RValue = NIL
--E 113
--S 114 of 235
)lisp (identity |$UserLevel|)
--R
--RValue = |development|
--E 114
--S 115 of 235

)lisp (identity |$whatOptions|)
--R
--IValue = (|operations| |categories| |domains| |packages| ...
--E 115
--S 116 of 235
)lisp (identity |$attributeDb|)
--R
--RValue = NIL
--E 116
--S 117 of 235
)lisp (identity $boot)
--R
--RValue = NIL
--E 117
--S 118 of 235
)lisp (identity |$cacheAlist|)
--R
--RValue = NIL
--E 118
--S 119 of 235
)lisp (identity |$cacheCount| )
--R
--RValue = 0
--E 119
--S 120 of 235
)lisp (identity |$CatOfCatDatabase|)
--R
--RValue = NIL
--E 120
--S 121 of 235
)lisp (identity |$CloseClient|)
--R
--RValue = 10
--E 121
--S 122 of 235
)lisp (identity |$coerceIntByMapCounter|)
--R
--RValue = 0
--E 122
--S 123 of 235
)lisp (identity |$compileMapFlag|)
--R

--RValue = NIL
--E 123
--S 124 of 235
)lisp (identity |$ConstructorCache|)
--R
--IValue = #<hash-table 08a68f18>
--E 124
--S 125 of 235
)lisp (identity |$constructors|)
--R
--RValue = NIL
--E 125
--S 126 of 235
)lisp (identity /countlist)
--R
--RValue = NIL
--E 126
--S 127 of 235
)lisp (identity $current-directory)
--R
--IValue = "/tmp/"
--E 127
--S 128 of 235
)lisp (identity |$currentFrameNum|)
--R
--RValue = 0
--E 128
--S 129 of 235
)lisp (identity |$currentLine|)
--R
--RValue = ")lisp (identity |$currentLine|)"
--E 129
--S 130 of 235
)lisp (identity $dalymode)
--R
--RValue = NIL
--E 130
--S 131 of 235
)lisp (identity |$defaultMsgDatabaseName|)
--R
--IValue = #p"/research/reference/mnt/ubuntu/doc/msgs/s2-us.msgs"
--E 131

--S 132 of 235
)lisp (identity |$dependeeClosureAlist|)
--R
--RValue = NIL
--E 132
--S 133 of 235
)lisp (identity $directory-list)
--R
--IValue = ("/research/reference/mnt/ubuntu/../../src/input/" ...
--E 133
--S 134 of 235
)lisp (identity |$displayStartMsgs| )
--R
--RValue = T
--E 134
--S 135 of 235
)lisp (identity |$domains|)
--R
--R
--R   >> System error:
--R   The variable |$domains| is unbound.
--R
--R   Continuing to read the file...
--R
--E 135
--S 136 of 235
)lisp (identity |$DomOfCatDatabase|)
--R
--RValue = NIL
--E 136
--S 137 of 235
)lisp (identity |$domainTraceNameAssoc|)
--R
--RValue = NIL
--E 137
--S 138 of 235
)lisp (identity |$doNotAddEmptyModeIfTrue|)
--R
--R
--R   >> System error:
--R   The variable |$doNotAddEmptyModeIfTrue| is unbound.
--R
--R   Continuing to read the file...

--R
--E 138
--S 139 of 235
)lisp (identity |$e|)
--R
--IValue = ((((|Category| (|modemap| (((|Category|) (|Category|)) (T *))))...
--E 139
--S 140 of 235
)lisp (identity |$echoLineStack|)
--R
--RValue = NIL
--E 140
--S 141 of 235
)lisp (identity /editfile)
--R
--IValue = #p"/tmp/u.input"
--E 141
--S 142 of 235
)lisp (identity |$EmptyEnvironment|)
--R
--RValue = ((NIL))
--E 142
--S 143 of 235
)lisp (identity |$env|)
--R
--RValue = ((NIL))
--E 143
--S 144 of 235
)lisp (identity *eof*)
--R
--RValue = NIL
--E 144
--S 145 of 235
)lisp (identity |$erMsgToss|)
--R
--RValue = NIL
--E 145
--S 146 of 235
)lisp (identity |$existingFiles|)
--R
--IValue = #<hash-table 08c5b230>
--E 146

--S 147 of 235
)lisp (identity |$fn|)
--R
--IValue = "/tmp/u.input"
--E 147
--S 148 of 235
)lisp (identity |$formulaOutputStream|)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 148
--S 149 of 235
)lisp (identity |$fortranOutputStream|)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 149
--S 150 of 235
)lisp (identity |$frameMessages|)
--R
--RValue = NIL
--E 150
--S 151 of 235
)lisp (identity |$frameRecord|)
--R
--RValue = NIL
--E 151
--S 152 of 235
)lisp (identity |$fromSpadTrace|)
--R
--RValue = NIL
--E 152
--S 153 of 235
)lisp (identity |$functionTable|)
--R
--RValue = NIL
--E 153
--S 154 of 235
)lisp (identity |$globalExposureGroupAlist|)
--R
--IValue = ((|basic| ...
--E 154
--S 155 of 235

)lisp (identity |$HistList|)
--R
--IValue = #0=(NIL NIL NIL NIL NIL NIL NIL NIL NIL NIL NIL NIL ...
--E 155
--S 156 of 235
)lisp (identity |$HistListAct|)
--R
--RValue = 0
--E 156
--S 157 of 235
)lisp (identity |$HistListLen|)
--R
--RValue = 20
--E 157
--S 158 of 235
)lisp (identity |$HistRecord|)
--R
--RValue = NIL
--E 158
--S 159 of 235
)lisp (identity |$inclAssertions|)
--R
--RValue = (AIX |CommonLisp|)
--E 159
--S 160 of 235
)lisp (identity |$InitialModemapFrame|)) )
--R
--RValue = ((NIL))
--E 160
--S 161 of 235
)lisp (identity in-stream)
--R
--RValue = #<synonym stream to *STANDARD-INPUT*>
--E 161
--S 162 of 235
)lisp (identity |$InteractiveMode|)
--R
--RValue = T
--E 162
--S 163 of 235
)lisp (identity |$InteractiveFrame| )
--R

--RValue = ((NIL))
--E 163
--S 164 of 235
)lisp (identity |$internalHistoryTable|)
--R
--RValue = NIL
--E 164
--S 165 of 235
)lisp (identity |$interpreterFrameName|)
--R
--IValue = |frame0|
--E 165
--S 166 of 235
)lisp (identity |$interpreterFrameRing|)
--R
--IValue = ((|frame0| ((NIL)) 1 T #0=(NIL NIL NIL NIL NIL ...
--E 166
--S 167 of 235
)lisp (identity |$intRestart|)
--R
--RValue = |restart|
--E 167
--S 168 of 235
)lisp (identity |$intTopLevel|)
--R
--RValue = |top_level|
--E 168
--S 169 of 235
)lisp (identity |$IOindex| )
--R
--RValue = 1
--E 169
--S 170 of 235
)lisp (identity |$JoinOfCatDatabase|)
--R
--RValue = NIL
--E 170
--S 171 of 235
)lisp (identity |$JoinOfDomDatabase|)
--R
--RValue = NIL
--E 171

--S 172 of 235
)lisp (identity |$lastPos|)
--R
--RValue = (|noposition|)
--E 172
--S 173 of 235
)lisp (identity |$lastUntraced|)
--R
--RValue = NIL
--E 173
--S 174 of 235
)lisp (identity |$letAssoc| )
--R
--RValue = NIL
--E 174
--S 175 of 235
)lisp (identity |$libQuiet|)
--R
--RValue = T
--E 175
--S 176 of 235
)lisp (identity $library-directory-list)
--R
--IValue = ("/research/reference/mnt/ubuntu/algebra/")
--E 176
--S 177 of 235
)lisp (identity |$localExposureData|)
--R
--IValue = #<vector 08cc0e8c>
--E 177
--S 178 of 235
)lisp (identity |$localExposureDataDefault|)
--R
--IValue = #<vector 08a687fc>
--E 178
--S 179 of 235
)lisp (identity |$lookupDefaults|)
--R
--R
--R   >> System error:
--R   The variable |$lookupDefaults| is unbound.
--R

--R   Continuing to read the file...
--R
--E 179
--S 180 of 235
)lisp (identity |$mathmlOutputStream|)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 180
--S 181 of 235
)lisp (identity |$mathTraceList|)
--R
--RValue = NIL
--E 181
--S 182 of 235
)lisp (identity |$mkTestInputStack|)
--R
--RValue = NIL
--E 182
--S 183 of 235
)lisp (identity |$msgAlist|)
--R
--RValue = NIL
--E 183
--S 184 of 235
)lisp (identity |$msgDatabase|)
--R
--RValue = NIL
--E 184
--S 185 of 235
)lisp (identity |$msgDatabaseName|)
--R
--RValue = NIL
--E 185
--S 186 of 235
)lisp (identity |$ncMsgList|)
--R
--RValue = NIL
--E 186
--S 187 of 235
)lisp (identity |$newConlist|)
--R
--RValue = NIL

--E 187
--S 188 of 235
)lisp (identity |$NonNullStream| )
--R
--RValue = "NonNullStream"
--E 188
--S 189 of 235
)lisp (identity |$nopos|)
--R
--RValue = (|noposition|)
--E 189
--S 190 of 235
)lisp (identity |$newcompErrorCount|)
--R
--RValue = 0
--E 190
--S 191 of 235
)lisp (identity nil)
--R
--RValue = NIL
--E 191
--S 192 of 235
)lisp (identity $newspad)
--R
--RValue = T
--E 192
--S 193 of 235
)lisp (identity |$NullStream|)
--R
--RValue = "NullStream"
--E 193
--S 194 of 235
)lisp (identity |$okToExecuteMachineCode|)
--R
--RValue = T
--E 194
--S 195 of 235
)lisp (identity |$openMathOutputStream|)
--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 195

--S 196 of 235
)lisp (identity |$operationNameList|)
--R
--RValue = NIL
--E 196
--S 197 of 235
)lisp (identity |$outputLibraryName|)
--R
--RValue = NIL
--E 197
--S 198 of 235
)lisp (identity |$OutputForm|)
--R
--RValue = (|OutputForm|)
--E 198
--S 199 of 235
)lisp (identity |$packages|)
--R
--R
--R   >> System error:
--R   The variable |$packages| is unbound.
--R
--R   Continuing to read the file...
--R
--E 199
--S 200 of 235
)lisp (identity /pretty)
--R
--RValue = NIL
--E 200
--S 201 of 235
)lisp (identity |$previousBindings|)
--R
--RValue = NIL
--E 201
--S 202 of 235
)lisp (identity |$PrintCompilerMessageIfTrue|)
--R
--RValue = NIL
--E 202
--S 203 of 235
)lisp (identity |$printLoadMsgs| )
--R

--S 204 of 235
)lisp (identity |$promptMsg|)
--R
--RValue = S2CTP023
--E 204
--S 205 of 235
)lisp (identity |$QuickLet|)
--R
--RValue = T
--E 205
--S 206 of 235
)lisp (identity |$quitTag|)
--R
--RValue = (NIL)
--E 206
--S 207 of 235
)lisp (identity $relative-directory-list)
--R
--IValue = ("/../../src/input/" "/doc/msgs/" "/../../src/algebra/" ...
--E 207
--S 208 of 235
)lisp (identity $relative-library-directory-list)
--R
--RValue = ("/algebra/")
--E 208
--S 209 of 235
)lisp (identity |$seen|)
--R
--R
--R   >> System error:
--R   The variable |$seen| is unbound.
--R
--R   Continuing to read the file...
--R
--E 209
--S 210 of 235
)lisp (identity |$SessionManager|)
--R
--RValue = 1
--E 210

--S 211 of 235
)lisp (identity |$setOptions|)
--R
--RValue = ((|breakmode| "execute break processing on error" |interpreter| LITERALS |$BreakMode| (|nobreak| |break| |query| |resume| |fastlinks| |quit|) |nobreak|) (|compiler| "Library compiler options" |interpreter| TREE |novar| ((|output| "library in which to place compiled code" |interpreter| FUNCTION |setOutputLibrary| NIL |htSetOutputLibrary|) (|input| "controls libraries from which to load compiled code" |interpreter| FUNCTION |setInputLibrary| NIL |htSetInputLibrary|))) (|debug| "debug options" |interpreter| TREE |novar| ((|lambdatype| "show type information for #1 syntax" |interpreter| LITERALS $LAMBDATYPE (|on| |off|) |off|) (|dalymode| "Interpret leading open paren as lisp" |interpreter| LITERALS $DALYMODE (|on| |off|) |off|))) (|expose| "control interpreter constructor exposure" |interpreter| FUNCTION |setExpose| NIL |htSetExpose|) (|functions| "some interpreter function options" |interpreter| TREE |novar| ((|cache| "number of function results to cache" |interpreter| FUNCTION |setFunctionsCache| NIL |htSetCache|) (|compile| "compile, don't just define function bodies" |interpreter| LITERALS |$compileDontDefineFunctions| (|on| |off|) |on|) (|recurrence| "specially compile recurrence relations" |interpreter| LITERALS |$compileRecurrence| (|on| |off|) |on|))) (|fortran| "view and set options for FORTRAN output" |interpreter| TREE |novar| ((|ints2floats| "where sensible, coerce integers to reals" |interpreter| LITERALS |$fortInts2Floats| (|on| |off|) |on|) (|fortindent| "the number of characters indented" |interpreter| INTEGER |$fortIndent| (0 NIL) 6) (|fortlength| "the number of characters on a line" |interpreter| INTEGER |$fortLength| (1 NIL) 72) (|typedecs| "print type and dimension lines" |interpreter| LITERALS |$printFortranDecs| (|on| |off|) |on|) (|defaulttype| "default generic type for FORTRAN object" |interpreter| LITERALS |$defaultFortranType| (REAL INTEGER COMPLEX LOGICAL CHARACTER) REAL) (|precision| "precision of generated FORTRAN objects" |interpreter| LITERALS |$fortranPrecision| (|single| |double|) |double|) (|intrinsic| "whether to use INTRINSIC FORTRAN functions" |interpreter| LITERALS |$useIntrinsicFunctions| (|on| |off|) |off|) (|explength| "character limit for FORTRAN expressions" |interpreter| INTEGER |$maximumFortranExpressionLength| (0 NIL) 1320) (|segment| "split long FORTRAN expressions" |interpreter| LITERALS |$fortranSegment| (|on| |off|) |on|) (|optlevel| "FORTRAN optimisation level" |interpreter| INTEGER |$fortranOptimizationLevel| (0 2) 0) (|startindex| "starting index for FORTRAN arrays" |interpreter| INTEGER |$fortranArrayStartingIndex| (0 1) 1) (|calling| "options for external FORTRAN calls" |interpreter| TREE |novar| ((|tempfile| "set location of temporary data files" |interpreter| FUNCTION |setFortTmpDir| (("enter directory name for which you have write-permission" DIRECTORY |$fortranTmpDir| |chkDirectory| "/tmp/")) NIL) (|directory| "set location of generated FORTRAN files" |interpreter| FUNCTION |setFortDir| (("enter directory name for which you have write-permission" DIRECTORY |$fortranDirectory| |chkDirectory| "./")) NIL) (|linker| "linker arguments (e.g. libraries to search)" |interpreter| FUNCTION |setLinkerArgs| (("enter linker arguments " STRING |$fortranLibraries| |chkDirectory| "-lxlf")) NIL))))) (|hyperdoc| "options in using HyperDoc" |interpreter| TREE |novar| ((|fullscreen| "use full screen for this facility" |interpreter| LITERALS |$fullScreenSysVars| (|on| |off|) |off|) (|mathwidth| "screen width for history output" |interpreter| INTEGER |$historyDisplayWidth| (0 NIL) 120))) (|help| "view and set some help options" |interpreter| TREE |novar| ((|fullscreen| "use fullscreen facility, if possible" |interpreter| LITERALS |$useFullScreenHelp| (|on| |off|) |on|))) (|history| "save workspace values in a history file" |interpreter| LITERALS |$HiFiAccess| (|on| |off|) |on|) (|messages| "show messages for various system features" |interpreter| TREE |novar| ((|any| "print the internal type of objects of domain Any" |interpreter| LITERALS |$printAnyIfTrue| (|on| |off|) |on|) (|autoload| "print file auto-load messages" |interpreter| LITERALS |$printLoadMsgs| (|on| |off|) |on|) (|bottomup| "display bottom up modemap selection" |development| LITERALS |$reportBottomUpFlag| (|on| |off|) |off|) (|coercion| "display datatype coercion messages" |development| LITERALS |$reportCoerceIfTrue| (|on| |off|) |off|) (|dropmap| "display old map defn when replaced" |interpreter| LITERALS |$displayDroppedMap| (|on| |off|) |off|) (|expose| "warning for unexposed functions" |interpreter| LITERALS |$giveExposureWarning| (|on| |off|) |off|) (|file| "print msgs also to SPADMSG LISTING" |development| LITERALS |$printMsgsToFile| (|on| |off|) |off|) (|frame| "display messages about frames" |interpreter| LITERALS |$frameMessages| (|on| |off|) |off|) (|highlighting| "use highlighting in system messages" |interpreter| LITERALS |$highlightAllowed| (|on| |off|) |off|) (|instant| "present instantiation summary" |development| LITERALS |$reportInstantiations| (|on| |off|) |off|) (|insteach| "present instantiation info" |development| LITERALS |$reportEachInstantiation| (|on| |off|) |off|) (|interponly| "say when function code is interpreted" |interpreter| LITERALS |$reportInterpOnly| (|on| |off|) |on|) (|naglink| "show NAGLink messages" |interpreter| LITERALS |$nagMessages| (|on| |off|) |on|) (|number| "display message number with message" |interpreter| LITERALS |$displayMsgNumber| (|on| |off|) |off|) (|prompt| "set type of input prompt to display" |interpreter| LITERALS |$inputPromptType| (|none| |frame| |plain| |step| |verbose|) |step|) (|selection| "display function selection msgs" |interpreter| LITERALS |$reportBottomUpFlag| (|on| |off|) |off|) (|set| "show )set setting after assignment" |interpreter| LITERALS |$displaySetValue| (|on| |off|) |off|) (|startup| "display messages on start-up" |interpreter| LITERALS |$displayStartMsgs| (|on| |off|) |on|) (|summary| "print statistics after computation" |interpreter| LITERALS |$printStatisticsSummaryIfTrue| (|on| |off|) |off|) (|testing| "print system testing header" |development| LITERALS |$testingSystem| (|on| |off|) |off|) (|time| "print timings after computation" |interpreter| LITERALS |$printTimeIfTrue| (|on| |off| |long|) |off|) (|type| "print type after computation" |interpreter| LITERALS |$printTypeIfTrue| (|on| |off|) |on|) (|void| "print Void value when it occurs" |interpreter| LITERALS |$printVoidIfTrue| (|on| |off|) |off|))) (|naglink| "options for NAGLink" |interpreter| TREE |novar| ((|host| "internet address of host for NAGLink" |interpreter| FUNCTION |setNagHost| (("enter host name" DIRECTORY |$nagHost| |chkDirectory| "localhost")) NIL) (|persistence| "number of (fortran) functions to remember" |interpreter| FUNCTION |setFortPers| (("Requested remote storage (for asps):" INTEGER |$fortPersistence| (0 NIL) 10)) NIL) (|messages| "show NAGLink messages" |interpreter| LITERALS |$nagMessages| (|on| |off|) |on|) (|double| "enforce DOUBLE PRECISION ASPs" |interpreter| LITERALS |$nagEnforceDouble| (|on| |off|) |on|))) (|output| "view and set some output options" |interpreter| TREE |novar| ((|abbreviate| "abbreviate type names" |interpreter| LITERALS |$abbreviateTypes| (|on| |off|) |off|) (|algebra| "display output in algebraic form" |interpreter| FUNCTION |setOutputAlgebra| (("display output in algebraic form" LITERALS |$algebraFormat| (|off| |on|) |on|) (BREAK $ALGEBRAFORMAT) ("where algebra printing goes (enter {em console} or a pathname)?" FILENAME |$algebraOutputFile| |chkOutputFileName| "console")) NIL) (|characters| "choose special output character set" |interpreter| FUNCTION |setOutputCharacters| NIL |htSetOutputCharacters|) (|fortran| "create output in FORTRAN format" |interpreter| FUNCTION |setOutputFortran| (("create output in FORTRAN format" LITERALS |$fortranFormat| (|off| |on|) |off|) (|break| |$fortranFormat|) ("where FORTRAN output goes (enter {em console} or a a pathname)" FILENAME |$fortranOutputFile| |chkOutputFileName| "console")) NIL) (|fraction| "how fractions are formatted" |interpreter| LITERALS |$fractionDisplayType| (|vertical| |horizontal|) |vertical|) (|html| "create output in HTML style" |interpreter| FUNCTION |setOutputHtml| (("create output in HTML format" LITERALS |$htmlFormat| (|off| |on|) |off|) (|break| |$htmlFormat|) ("where HTML output goes (enter {em console} or a pathname)" FILENAME |$htmlOutputFile| |chkOutputFileName| "console")) NIL) (|length| "line length of output displays" |interpreter| INTEGER $LINELENGTH (10 245) 77) (|mathml| "create output in MathML style" |interpreter| FUNCTION |setOutputMathml| (("create output in MathML format" LITERALS |$mathmlFormat| (|off| |on|) |off|) (|break| |$mathmlFormat|) ("where MathML output goes (enter {em console} or a pathname)" FILENAME |$mathmlOutputFile| |chkOutputFileName| "console")) NIL) (|openmath| "create output in OpenMath style" |interpreter| FUNCTION |setOutputOpenMath| (("create output in OpenMath format" LITERALS |$openMathFormat| (|off| |on|) |off|) (|break| |$openMathFormat|) ("where TeX output goes (enter {em console} or a pathname)" FILENAME |$openMathOutputFile| |chkOutputFileName| "console")) NIL) (|script| "display output in SCRIPT formula format" |interpreter| FUNCTION |setOutputFormula| (("display output in SCRIPT format" LITERALS |$formulaFormat| (|off| |on|) |off|) (|break| |$formulaFormat|) ("where script output goes (enter {em console} or a a pathname)" FILENAME |$formulaOutputFile| |chkOutputFileName| "console")) NIL) (|scripts| "show subscripts,... linearly" |interpreter| LITERALS |$linearFormatScripts| (|on| |off|) |off|) (|showeditor| "view output of )show in editor" |interpreter| LITERALS |$useEditorForShowOutput| (|on| |off|) |off|) (|tex| "create output in TeX style" |interpreter| FUNCTION |setOutputTex| (("create output in TeX format" LITERALS |$texFormat| (|off| |on|) |off|) (|break| |$texFormat|) ("where TeX output goes (enter {em console} or a pathname)" FILENAME |$texOutputFile| |chkOutputFileName| "console")) NIL))) (|quit| "protected or unprotected quit" |interpreter| LITERALS |$quitCommandType| (|protected| |unprotected|) |protected|) (|streams| "set some options for working with streams" |interpreter| TREE |novar| ((|calculate| "specify number of elements to calculate" |interpreter| FUNCTION |setStreamsCalculate| (("number of initial stream elements you want calculated" INTEGER |$streamCount| (0 NIL) 10)) NIL) (|showall| "display all stream elements computed" |interpreter| LITERALS |$streamsShowAll| (|on| |off|) |off|))) (|system| "set some system development variables" |development| TREE |novar| ((|functioncode| "show gen. LISP for functions when compiled" |development| LITERALS |$reportCompilation| (|on| |off|) |off|) (|optimization| "show optimized LISP code" |development| LITERALS |$reportOptimization| (|on| |off|) |off|) (|prettyprint| "prettyprint BOOT func's as they compile" |development| LITERALS $PRETTYPRINT (|on| |off|) |on|))) (|userlevel| "operation access level of system user" |interpreter| LITERALS |$UserLevel| (|interpreter| |compiler| |development|) |development|))
--E 211
--S 212 of 235 obsolete |$shoeReadLineFunction|
)lisp nil
--R
--RValue = NIL
--E 212
--S 213 of 235
)lisp (identity |$slamFlag|)
--R
--RValue = NIL
--E 213
--S 214 of 235
)lisp (identity /sourcefiles)
--R
--RValue = NIL
--E 214
--S 215 of 235
)lisp (identity |$sourceFiles|)
--R
--RValue = NIL
--E 215
--S 216 of 235
)lisp (identity /spacelist)
--R
--RValue = NIL
--E 216
--S 217 of 235
)lisp (identity $spad)
--R
--RValue = T
--E 217
--S 218 of 235
)lisp (identity $spadroot)
--R
--IValue = "/research/reference/mnt/ubuntu"
--E 218
--S 219 of 235
)lisp (identity |$texOutputStream|)

--R
--RValue = #<synonym stream to *TERMINAL-IO*>
--E 219
--S 220 of 235
)lisp (identity /timerlist)
--R
--RValue = NIL
--E 220
--S 221 of 235
)lisp (identity |$timerTicksPerSecond|)
--R
--RValue = 100
--E 221
--S 222 of 235
)lisp (identity |$tracedMapSignatures|)
--R
--RValue = NIL
--E 222
--S 223 of 235
)lisp (identity |$tracedModemap|)
--R
--RValue = NIL
--E 223
--S 224 of 235
)lisp (identity |$tracedSpadModemap|)
--R
--RValue = NIL
--E 224
--S 225 of 235
)lisp (identity |$traceErrorStack|)
--R
--R
--R   >> System error:
--R   The variable |$traceErrorStack| is unbound.
--R
--R   Continuing to read the file...
--R
--E 225
--S 226 of 235
)lisp (identity $traceletflag)
--R
--RValue = NIL
--E 226

--S 227 of 235
)lisp (identity |$traceletFunctions|)
--R
--RValue = NIL
--E 227
--S 228 of 235
)lisp (identity |$undoFlag|)
--R
--RValue = T
--E 228
--S 229 of 235
)lisp (identity nil)
--R
--RValue = NIL
--E 229
--S 230 of 235
)lisp (identity |$UserAbbreviationsAlist|)
--R
--RValue = NIL
--E 230
--S 231 of 235
)lisp (identity |$variableNumberAlist|)
--R
--RValue = NIL
--E 231
--S 232 of 235
)lisp (identity |$Void|)
--R
--RValue = (|Void|)
--E 232
--S 233 of 235
)lisp (identity |$writifyComplained|)
--R
--R
--R   >> System error:
--R   The variable |$writifyComplained| is unbound.
--R
--R   Continuing to read the file...
--R
--E 233
--S 234 of 235
)lisp (identity /wsname)

--R
--RValue = NOBOOT
--E 234
--S 235 of 235
)lisp (identity |$xdatabase|)
--R
--RValue = NIL
--E 235

)spool
)lisp (bye)

———-

References^{[1] nothing}

33